OpenStudy (drewkatski):

Hey, can someone please explain how to do these questions? I really am lost :x If f (x)= 1-2x Find: a) f (5) b) f (-5) c) f (1/4) Thankyou in advance

OpenStudy (anonymous):

f(x)=1-2x f (5)=1-2(5) = 1-10 =-9

OpenStudy (anonymous):

can you find f(-5)?

OpenStudy (drewkatski):

ohh ok, i see... f (-5) = 1-2(-5) = 1--10 = 11 like that right?

OpenStudy (drewkatski):

@mustafa2014

OpenStudy (anonymous):

right :)

OpenStudy (drewkatski):

:D thanks

OpenStudy (anonymous):

Yw

OpenStudy (drewkatski):

There is one other similar one that i dont understand too If f (x) = 3x^2-4

OpenStudy (drewkatski):

Write f (2x) in its simplest form

OpenStudy (anonymous):

f (2x)= 3(2x)^2-4 =3(4x^2)-4=12x^2-4